OpenStudy (anonymous):

integrate x/(x^2+1)dx ?

OpenStudy (anonymous):

try \[u=x^2+1\] \[du=2xdx\] \[\frac{1}{2}du=dx\] then \[\frac{1}{2}\int \frac{1}{u}du\] get \[\frac{1}{2}\ln(u)\]

OpenStudy (anonymous):